OpenStudy (anonymous):

4cos5theta-2sqrt(3)=0

OpenStudy (anonymous):

\[4\cos(5 \theta)-2 \sqrt{3}=0\]

OpenStudy (anonymous):

\[2\cos(5 \theta)= \sqrt{3}\]

OpenStudy (anonymous):

\[\cos(5 \theta) = \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

\[5 \theta = \cos^{-1}(\frac{ \sqrt{3} }{ 2 })\]

OpenStudy (anonymous):

\[\theta = \frac{ 1 }{ 5 }\cos^{-1}(\frac{ \sqrt{3} }{ 2 })\]

OpenStudy (anonymous):

ok I need to give all solutions

OpenStudy (anonymous):

arccos(sqrt3/2)=30 degree or pi/6

OpenStudy (anonymous):

then multiply by 1/5

OpenStudy (anonymous):

well as you can see that you have a positive angle, which means you have to calculate the angles in the first and fourth quarter, and add 5x theta round

OpenStudy (anonymous):

alright thanks! I got it. one quick question: would the problem 4cos^2=3=0 change any since there is no theta?

OpenStudy (anonymous):

yes the answer would be different of you square the cosine

OpenStudy (anonymous):

there would also be more solutions correct?

OpenStudy (anonymous):

if*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok and for sin 1/3t=1/2 would one solution be pi/2?

OpenStudy (anonymous):

sorry back

OpenStudy (anonymous):

\[\sin(\frac{ 1 }{ 3t })=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 3t } = \sin^{-1}(0.5)\]

OpenStudy (anonymous):

\[t = \frac{ 3 }{ \arcsin(0.5) }\]

OpenStudy (anonymous):

\[t = 3arccsc(0.5)\]

OpenStudy (anonymous):

i think one of the purposes of this question is to teach you the difference between inverse of a function and reciprocal